# Large Scale White Noise and Cosmology

### v1

@@ V.1 Cosmological Inhomogeneities
Cosmological inhomogeneities can be described as inhomogeneous metric and matter perturbations relative to a homogeneous and isotropic Friedman-Lemaitre-Robertson-Walker (FLRW) background. These perturbations are described as a realization of a stochastic process which is itself homogeneous and isotropic, consistent with the symmetries of the underlying space-time. FLRW space-times can be finite or infinite in spatial extent, but are empirically constrained to be sufficiently large so as to be modeled as infinite. The spatial sections may be positively curved, negatively curved, or flat (zero curvature), but the curvature is empirically constrained to be sufficiently small so as to be modeled as zero. Here, we assume the inhomogeneities are spatially homogeneous and isotropic in flat, infinite Euclidean spatial sections. The LSWN formalism can be generalized to non-flat or non-infinite FLRW space-times, just not here.
Unlike classical fluid dynamics, which operates under action-at-a-distance Newtonian gravity, cosmological inhomogeneities are governed by local equations (PDEs). Like classical fluid dynamics, the equations of motion for density, pressure, etc are non-linear with quadratic leading order non-linearities, just as we have found in the Newtonian cosmology illustrated in §IV. Our universe is inferred to be nearly homogeneous at early times and on large scales, so the small inhomogeneities are usually modeled as a perturbative expansion about homogeneity. Usually only the first order (linear theory) dynamics is considered. Pseudo-linear systems like this have been the focus of this paper so far. As with the illustrations in §IV, one should expect that the leading order Born approximation (LOBA) developed previously should provide an accurate description of the effect of non-linearities.
@@ V.3 Curvature Inhomogeneities
$\Delta\rho\equiv\rho-\frac{1}{27\pi\,G}\,\theta^{2}$ (172)
@@
Because $\Delta\rho$, unlike $\mathcal{R}$, is a locally measurable quantity, and because physical interactions are local, the evolution of $\Delta\rho$ must be described by local equations, i.e. by PDEs. In GR, energy-momentum is locally but not globally conserved, so as expected $\Delta\rho$ does not give rise to any conserved current [6]. Other quantities may be exactly or approximately conserved in cosmology, but it is not these quantities which couple linearly with the gravitational field. For example, entropy may be conserved during the radiation era, but the entropy density is related to the mass density by a non-linear relation $\rho\propto s^{4/3}$. Thus, while the power spectrum of the conserved $s$ may not develop LSWN, the power spectrum of the non-conserved $\rho$ must. [^18] There is the more subtle issue of which spatial coordinates are appropriate to describe the observations of the power spectrum. Non-linearly the large scale power spectrum will depend on the coordinates e.g. Lagrangian vs. Eulerian, and may have LSWN in one but not the other, even though the two coincide in linear theory.
@@ Appendix C LSWN Kernel For Self-similar Local Systems
In §II.4 & LABEL:sec:AsymptotesAmplitudesPhases we define the functions ${\cal M}_{\pm}[\varphi]$ and ${\cal M}_{\pm}[\varphi]={\cal M}_{+}[\varphi]+{\cal M}_{-}[\varphi]$ which give the LSWN for $n=p=2$ self similar systems. These are defined the integrals of eq. 87 using eq.s 73 & 68. These integrals can be performed analytically:

### v2

@@ V.1 Cosmological Inhomogeneities
Cosmological inhomogeneities can be described as inhomogeneous metric and matter perturbations relative to a homogeneous and isotropic Friedman-Lemaitre-Robertson-Walker (FLRW) background. These perturbations are described as a realization of a stochastic process which is itself homogeneous and isotropic, consistent with the symmetries of the underlying space-time. FLRW space-times can be finite or infinite in spatial extent, but are empirically constrained to be sufficiently large so as to be modeled as infinite. The spatial sections may be positively curved, negatively curved, or flat (zero curvature), but the curvature is empirically constrained to be sufficiently small that one may only consider perturbations about a flat cosmology. Here, we assume the inhomogeneities are spatially homogeneous and isotropic in flat, infinite Euclidean spatial sections. The LSWN formalism can be generalized to non-flat or non-infinite FLRW space-times, just not here.
Unlike classical fluid dynamics, which operates under action-at-a-distance Newtonian gravity, cosmological inhomogeneities are governed by local equations (PDEs). Like classical fluid dynamics, the equations of motion for density, pressure, etc. are non-linear with quadratic leading order non-linearities, just as we have found in the Newtonian cosmology illustrated in §IV. Our universe is inferred to be nearly homogeneous at early times and on large scales, so the small inhomogeneities are usually modeled as a perturbative expansion about homogeneity. Usually only the first order (linear theory) dynamics is considered. Pseudo-linear systems like this have been the focus of this paper so far. As with the illustrations in §IV, one should expect that the leading order Born approximation (LOBA) developed previously should provide an accurate description of the effect of non-linearities.
@@ V.3 Curvature Inhomogeneities
$\Delta\rho\equiv\rho-\frac{1}{24\pi\,G}\,\theta^{2}$ (173)
@@
Because $\Delta\rho$, unlike $\mathcal{R}$, is a locally measurable quantity, and because physical interactions are local, the evolution of $\Delta\rho$ must be described by local equations, i.e. by PDEs. In GR, energy-momentum is locally but not globally conserved, so as expected $\Delta\rho$ does not give rise to any conserved current [6]. Other quantities may be exactly or approximately conserved in cosmology, but it is not these quantities which couple linearly with the gravitational field. For example, entropy may be conserved during the radiation era, but the fluid entropy density is related to the mass density by a non-linear relation $\rho\propto s^{4/3}$. Furthermore, even if entropy is conserved it will still develop LSWN. As pointed out in §III.3 in order or $s$ to avoid developing LSWN it must not only be conserved but also have pure growing mode and sub-Poissonian initial conditions. This cannot happen if for no other reason that, unlike $\Delta\rho$, $s$ is not a perturbation variable, i.e. $s=0$ does not describe the unperturbed cosmology. Instead $s$ will developed LSWN related in the usual way to the relic curvature LSWN described below.
@@ V.5.4 Heretical Models
V.6 Simplicity
Explicit representation of Einstein equations are long and complex expressions, even in the case of planar symmetry. They are non-linear to arbitrary order. The procedure used here, which we claim is accurate in the pseudo-linear regime, greatly reduces the relevant complexity.
Scalar curvature perturbations are described by a 2nd order, $p=2$, PDE at linear order. It is clear, even from classical hydrodynamics, that the leading order non-linearities are quadratic, $n=2$. We claim that the leading order Born approximation (LOBA) is accurate in the pseudo-linear regime. LOBA corrects the the linear solutions to 2nd order using the linear solutions. These require initial conditions.
Determining appropriate initial conditions is physics and may be the hardest part of the problem. Generally one would idealize by specifying initial conditions at the initial singularity. One would probably specify a guiding variable (see III.3). This must be a locally measurable covariant quantity e.g. $\Delta\rho$ or perhaps $\Delta\equiv\Delta\rho/\rho$ or perhaps $\delta\Omega\equiv\Delta\rho/(\rho-\Delta\rho)$. They all differ.
Since the 2nd order PDE must be spatially homogeneous and isotropic this reduces the complicated Einstein equations to an EoM requiring only a few temporal functions (see §II.3). Furthermore in an idealized radiation era which is self-similar because of the constant sound speed these functions reduce to only seven numbers $\zeta_{1}$-$\zeta_{7}$ to describe the dynamics. It is likely in the radiation era that non-linearities cease to be active early on and mostly manifest as a relic change of the spectrum which then evolves linearly (see §III.1). This relic depends on the initial conditions, two numbers, $\mu$ and $\nu$, which describe the linear evolution plus one parameter, $\mathfrak{W}_{+}$, which describes the non-linearly induced LSWN. We know how to model linear evolution so we know $\mu$ and $\nu$. Thus GR non-linear evolution of curvature inhomogeneities in the radiation era reduce to one number $\mathfrak{W}_{+}$. The radiation era relic spectrum can be used as early universe initial conditions for later evolution. In some sense we have reduced the important GR non-linear dynamics to one unknown number without specifying any of the details of GR. This number only serves to normalize the correction to the early radiation era power spectrum.
@@ Appendix C LSWN Kernel For Self-similar Local Systems
In §II.4, we define the functions ${\cal M}_{\pm}[\varphi]$ and ${\cal M}_{\pm}[\varphi]={\cal M}_{+}[\varphi]+{\cal M}_{-}[\varphi]$ which give the LSWN for $n=p=2$ self similar systems. These are defined the integrals of eq. 87 using eq.s 73 & 68. These integrals can be performed analytically: